2.8-3x=20 what is x?

Answer by checkley77(12569)

2.8-3x=20
-3x=20-2.8
-3x=17.2
x=17.2/-3
x=-17.2/3 ans.
Proof:
2.8-3*-17.2/3=20
2.8+17.2=20
20=20